The line (x+6)/5=(y+10)/3=(z+14)/8 is the hypotenuse of a isosceles right angled triangle whose opposite vertex is (7,2,4). Find the equation of the remaining sides

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
I am giving you a hint.
The required lines are the lines passing through (7,2,4) which intersect the given line at an angle of 45°.
Lines will be
x-7/3 = y-2/6 = z-2/4
And
x-7/2 = y -2/-3 = z-4/6
 
Regards
Arun (askIITians forum expert)